# Optimal Control of Nonlinear Systems with Unknown Dynamics

### v2

@@ 1 Introduction
Optimal control theory provides a foundational mathematical framework for the design of control strategies aimed at minimizing user-defined cost functions, typically under the assumption of known system dynamics [1]. For systems with unknown dynamics, model-free methods have emerged as a promising alternative for deriving the closed-loop optimal controller without using the system dynamics [2]. These approaches commonly assume unknown dynamics as a Markov decision process and use parameterized policies as closed-loop controllers. They are often implemented using an actor-critic structure [3], where the critic evaluates policy performance using observed data, and the actor updates the policy parameters based on the critic’s feedback. Here, the critic approximates the expectation of the cost function by minimizing the temporal difference error [4] without using the system dynamics, while the actor optimizes the policy by minimizing the critic, which is also referred to as policy gradient methods [5]. Notable contributions in this domain include the deep Q network [6], and proximal policy optimization [7], which are well-suited for discrete action spaces. For continuous action spaces, deterministic policy gradient algorithms [8] have been incorporated into the deep Q network framework [9]. Nevertheless, actor-critic methods typically require extensive trial-and-error iterations to identify the closed-loop optimal controllers [10], which can present challenges in real-world applications.
@@
To improve the efficiency of model-free methods in identifying closed-loop optimal controllers, recent attention has been given to data-driven model-based methods, which involve estimating system dynamics and subsequently deriving closed-loop optimal controllers using the learned models. For example, integral reinforcement learning [11] directly determines optimal feedback gains from input-output data collected along system trajectories, typically assuming linear system dynamics. For systems with completely unknown nonlinear dynamics, several studies [12, 13, 14, 15, 16] focus on learning unknown dynamics using neural networks, followed by leveraging classical model-based control strategies such as model predictive control (MPC) to compute optimal control actions. In the area of model-based reinforcement learning [17], popular methods [18, 19] leverage Gaussian processes to approximate system dynamics, while works [20, 21] utilize deep neural networks (DNNs) to learn the dynamics. These techniques facilitate the generation of closed-loop optimal controllers by performing policy gradient updates using the learned models to directly minimize the user-defined cost function. Despite the successful integration of dynamic learning and policy design, several challenges persist, particularly due to the nonlinearity of learned dynamic models. These challenges include difficulties in verifying critical properties of the learned dynamics, such as controllability and stability, as noted in [22]. Furthermore, the significant nonlinearity in the learned dynamics often results in substantial computational complexity, which complicates practical implementation [23].
@@
The Koopman operator [24] offers an alternative approach to approximate nonlinear systems with linear dynamics based on state-control pairs [25, 26, 27], enabling the evaluation of key properties of the learned dynamics, such as observability and controllability. Techniques like extended dynamic mode decomposition [28] transform the state space into a higher-dimensional space, making the dynamics approximately linear through carefully selected lifting functions. To address the complexity of choosing these functions, recent studies, for instance, [29], has employed deep learning techniques to discover the eigenfunctions of the Koopman operator. Deep Koopman operator (DKO) methods [30, 31, 32, 33] utilize DNNs as lifting functions, optimizing them with state-control pairs and a properly defined loss function. Model-based controllers, such as MPC, can then be integrated with the learned Koopman dynamics model [30]. Nevertheless, inaccuracies in the learned model may result in cumulative errors during cost function computation when propagating system states, thereby affecting the reliability of the approach.
@@
Motivated by the above, this paper proposes a data-driven framework to derive the closed-loop optimal controller leveraging the combined benefits of actor-critic algorithms and dynamics learning using the DKO approach. This is achieved by decomposing the bilevel optimization problem of the actor-critic structure through the integration of the DKO into the framework, enabling independent tuning of the dynamics approximator and the critic. The primary contribution of this paper lies in developing a parameter-tuning framework that simultaneously approximates unknown dynamics, refines the critic, and optimizes the parameterized policy based on the critic’s evaluation using a one-time-step prediction from the DKO. The proposed approach is distinct from recent advancements in integrating the Koopman operator into reinforcement learning (RL). For instance, [34] employs the Koopman operator to augment the static offline datasets during training, whereas the present work focuses on estimating the policy gradient using the Koopman operator. Similarly, [35] leverages the Koopman operator to directly approximate the critic, enhancing the actor-critic framework, while the proposed method primarily leverages the DKO to predict future system states, facilitating improved policy optimization.
@@
[…]
2 Problem
@@
$\boldsymbol{x}(t+1)=\boldsymbol{f}(\boldsymbol{x}(t),\boldsymbol{u}(t)),\quad\boldsymbol{x}(0)=\boldsymbol{x}_{0},$ (1)
where $t=0,1,2,\cdots$ denotes the time index, $\boldsymbol{x}(t)\in\mathcal{X}\subset\mathbb{R}^{n}$ and $\boldsymbol{u}(t)\in\mathcal{U}\subset\mathbb{R}^{m}$ denote the system state and control input at time $t$, respectively, $\boldsymbol{x}_{0}$ is the given arbitrary vector, and $\boldsymbol{f}:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$ denotes the time-invariant, Lipschitz continuous, and unknown dynamics mapping. We assume $\mathcal{X}$ is a countably infinite state space. Notably, for the remainder of this paper, we distinguish between system state-input variables and fixed observed state-input pairs by denoting $\boldsymbol{x}_{t}$ and $\boldsymbol{u}_{t}$ as constant state and input vectors observed at time $t$, respectively.
@@
$J_{\mu,t^{*}}=\sum_{t=t^{*}}^{\infty}\gamma^{t-t^{*}}c(\boldsymbol{x}(t),\boldsymbol{u}(t)),$ (2)
where $c:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$ denotes a bounded and Lipschitz continuous stage cost function, $10<\gamma<$ 1 is a discount factor, and the control inputs are assumed to be chosen from a policy obeying the following form:
@@
Assume that the closed-loop optimal controller that minimizes $J_{\mu,t^{*}}$ in (2) can be represented by $\boldsymbol{\mu}(\boldsymbol{x}(t),\boldsymbol{\theta}^{\mu*})$. The problem of interest is to find a data-driven framework to achieve $\boldsymbol{\theta}^{\mu*}$. Correspondingly, given any arbitrary $\boldsymbol{x}(t^{*})=\boldsymbol{x}_{t^{*}}$, this paper aims to solve the following optimization problem without knowing the actual dynamics $\boldsymbol{f}$ in (1):
[…]
@@
The optimization problem in (4) is a well-established optimal control problem when the dynamics $\boldsymbol{f}$ in (1) is known [36]. In the case where $\boldsymbol{f}$ is unknown, one approach to solve (4) involves model-free methods, such as reinforcement learning techniques, which typically require a substantial amount of data to achieve an optimal solution. In the following section, we will propose a framework that concurrently approximates $\boldsymbol{f}$ using the Koopman operator and optimizes $\boldsymbol{\theta}^{\mu}$ to find the $\boldsymbol{\theta}^{\mu*}$. The method is shown to have better efficiency of computing $\boldsymbol{\theta}^{\mu*}$ in later simulations.
@@ 3.1 Challenges and Key Ideas
An iterative method to find $\boldsymbol{\theta}^{\mu*}$ in (4) is to employ the gradient descent method. Specifically, let $k=0,1,2,\cdots$ denote the iteration index, $\boldsymbol{\theta}_{k}^{\mu}\in\mathbb{R}^{q}$ denote the estimation of $\boldsymbol{\theta}^{\mu*}$ at the $k$-th iteration, and $\alpha_{k}^{\mu}$ represent the step size at the $k$-th iteration corresponding to $\boldsymbol{\theta}_{k}^{\mu}$. Given an arbitrary initial $\boldsymbol{\theta}_{0}^{\mu}$, the parameter $\boldsymbol{\theta}_{k}^{\mu}$ is updated iteratively using the gradient of $J_{\mu,t^{*}}$ as follows:
$\boldsymbol{\theta}_{k+1}^{\mu}=\boldsymbol{\theta}_{k}^{\mu}-\alpha_{k}^{\mu}\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}(\boldsymbol{\theta}_{k}^{\mu}),\quad\boldsymbol{\theta}_{0}^{\mu}\ \text{given},$ (5)
@@
$\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}=$ $\nabla_{\boldsymbol{\theta}^{\mu}}c(\boldsymbol{x}(t^{*}),\boldsymbol{u}(t^{*}))+\gamma\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}+1},$ (6)
$=$ $\frac{\partial c_{t^{*}}}{\partial\boldsymbol{u}_{t^{*}}}\frac{\partial\boldsymbol{u}_{t^{*}}}{\partial\boldsymbol{\theta}^{\mu}}+\sum_{t=t^{*}+1}^{\infty}\gamma^{t-t^{*}}(\frac{\partial c_{t}}{\partial\boldsymbol{u}(t)}\frac{\partial\boldsymbol{u}(t)}{\partial\boldsymbol{\theta}^{\mu}}$
@@
Here, $c_{t}\coloneqq c(\boldsymbol{x}(t),\boldsymbol{u}(t))$ is introduced for notation brevity, and the selection of $\alpha_{k}^{\mu}$ will be discussed in detail later in this paper. Throughout this paper, we refer to $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}$ as the policy gradient.
Two key challenges arise when computing the policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}$ in (6) due to the unknown dynamics $f$ in (1). First, the system dynamics $\boldsymbol{f}$ is assumed to be unknown in the present work, making the gradient $\frac{\partial\boldsymbol{x}(t)}{\partial\boldsymbol{u}(t-1)}$ unknown as well. Second, calculating $\sum_{t=t^{*}+1}^{\infty}\frac{\partial c_{t}}{\partial\boldsymbol{x}(t)}\frac{\partial\boldsymbol{x}(t)}{\partial\boldsymbol{u}(t-1)}\frac{\partial\boldsymbol{u}(t-1)}{\partial\boldsymbol{\theta}^{\mu}}+\frac{\partial c_{t}}{\partial\boldsymbol{u}(t)}\frac{\partial\boldsymbol{u}(t)}{\partial\boldsymbol{\theta}^{\mu}}$ requires future system states $\boldsymbol{x}(t^{*}+1),\boldsymbol{x}(t^{*}+2),\cdots$ that evolve according to system dynamics. Even if one can approximate the dynamics $\boldsymbol{f}$ with small estimation errors, these errors may still accumulate over an infinite time horizon, potentially resulting in inaccuracies in calculating $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}$.
To address the two challenges in computing the policy gradient, we propose a framework that applies the deep Koopman operator (DKO) to approximate the unknown dynamics $\boldsymbol{f}$ in a linear form, enabling efficient approximation of $\frac{\partial\boldsymbol{x}(t+1)}{\partial\boldsymbol{u}(t)}$ and potentially improving gradient-based optimization. Subsequently, we use the temporal difference error technique [37] to approximate $J_{\mu,t^{*}+1}$ such that $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}+1}$ can be estimated using only a one-time-step prediction from the DKO to reduce the computational complexity.

### v3

@@ 1 Introduction
Optimal control theory provides a foundational mathematical framework for the design of control strategies aimed at minimizing user-defined cost functions, typically under the assumption of known system dynamics [anderson2007optimal]. For systems with unknown dynamics, model-free reinforcement learning (RL) [sutton2018reinforcement] has emerged as a promising alternative to derive closed-loop optimal controllers directly from data. In RL, the unknown dynamics are modeled as a Markov decision process, and parameterized policies are employed as closed-loop controllers. A widely adopted structure is the actor–critic framework [konda1999actor], where the critic evaluates the policy using observed data and the actor updates the policy parameters based on the critic’s feedback. The critic approximates the expected cost by minimizing the temporal-difference error [sutton1988learning] without requiring explicit knowledge of the system dynamics, while the actor improves the policy by optimizing against the critic. Prominent advances in RL include deep Q networks [mnih2015human], proximal policy optimization [schulman2017proximal], and deterministic policy gradient algorithms [lillicrap2015continuous], which extend Q-learning ideas to continuous action spaces [silver2014deterministic]. Despite these successes, RL methods typically demand a large number of trial-and-error interactions to identify near-optimal controllers [xiong2022deterministic], posing significant challenges for real-world deployment.
@@
To improve the efficiency of model-free methods in identifying closed-loop optimal controllers, recent attention has been given to data-driven model-based methods, which involve estimating system dynamics and subsequently deriving closed-loop optimal controllers using the learned models. For example, integral RL [lewis2012optimal] directly determines optimal feedback gains from input-output data collected along system trajectories, typically assuming linear system dynamics. For systems with completely unknown nonlinear dynamics, several studies [hernandaz1990neural, miller1990real, draeger1995model, chua2018deep, askari2025model] focus on learning unknown dynamics using neural networks, followed by leveraging classical model-based control strategies such as model predictive control (MPC) to compute optimal control actions. In the area of model-based RL [doya2000reinforcement], popular methods [deisenroth2011pilco, chebotar2017combining] leverage Gaussian processes to approximate system dynamics, while works [deisenroth2011learning, levine2016end] utilize deep neural networks (DNNs) to learn the dynamics. These techniques facilitate the generation of closed-loop optimal controllers by performing policy gradient updates using the learned models to directly minimize the user-defined cost function. Despite the successful integration of dynamic learning and policy design, several challenges persist, particularly due to the nonlinearity of learned dynamic models. These challenges include difficulties in verifying critical properties of the learned dynamics, such as controllability and stability, as noted in [tanaka1996approach]. Furthermore, the significant nonlinearity in the learned dynamics often results in substantial computational complexity, which complicates practical implementation [polydoros2017survey].
@@
The Koopman operator [proctor2016dynamic] offers an alternative approach to approximate nonlinear systems with linear dynamics based on state-control pairs [mezic2015applications, proctor2018generalizing, mauroy2016linear], enabling the evaluation of key properties of the learned dynamics, such as observability and controllability. Techniques like extended dynamic mode decomposition [korda2018convergence] transform the state space into a higher-dimensional space, making the dynamics approximately linear through carefully selected lifting functions. Operator-theoretic methods have further applied Koopman-based lifting of Hamiltonian systems via the Pontryagin maximum principle [villanueva2021towards]. To address the complexity of choosing lifting functions, recent studies, for instance, [lusch2017data] has employed deep learning techniques to discover the eigenfunctions of the Koopman operator. Deep Koopman operator (DKO) methods [korda2018linear, dk, morton2018deep, hao2024deep] utilize DNNs as lifting functions, optimizing them with state-control pairs and a properly defined loss function. Model-based controllers, such as MPC, can then be integrated with the learned Koopman dynamics model [korda2018linear]. Nevertheless, inaccuracies in the learned model may result in cumulative errors during cost function computation when propagating system states, thereby affecting the reliability of the approach.
@@
Motivated by the above, this paper proposes a data-driven framework to derive the closed-loop optimal controller via the combined benefits of actor-critic concepts and dynamics learning using the DKO approach. This is achieved by decomposing the bilevel optimization problem of the actor-critic structure through the integration of the DKO, enabling independent tuning of the dynamics approximator and the critic. The proposed approach is distinct from recent advancements in integrating the Koopman operator into RL. For instance, [weissenbacher2022koopman] employs the Koopman operator to augment the static offline datasets during training, whereas the present work focuses on estimating the policy gradient using the Koopman operator. Similarly, [rozwood2024koopman] leverages the Koopman operator to directly approximate the critic, enhancing the actor-critic framework, while the proposed method primarily leverages the DKO to predict future system states, facilitating improved policy optimization. The main contributions are summarized as follows:
  • To accelerate dynamics learning, the dynamics identification problem is formulated as a four-variable optimization problem. A multivariable update rule grounded in DKO is developed, and its convergence is analyzed. The results demonstrate faster convergence compared with standard single-parameter updates.
  • To mitigate model error accumulation during state propagation when computing the infinite-horizon cost gradient with respect to policy parameters, a DKO-based policy gradient estimation scheme is introduced. By embedding DKO learning into the actor–critic framework, the proposed method enables concurrent dynamics learning, critic refinement, and policy optimization using only one-step predictions from the DKO dynamics.
@@
Notations. Let $\parallel\cdot\parallel$ be the Euclidean norm. For a matrix $A\in\mathbb{R}^{n\times m}$, $\parallel A\parallel_{F}$ denotes its Frobenius norm, $A^{\prime}$ denotes its transpose, $A^{\dagger}$ denotes its Moore-Penrose pseudoinverse, $Tr(A)$ denotes its trace, and $\lambda_{\mathrm{min}}(AA^{\prime})$ is the minimum eigenvalue of $AA^{\prime}$. $\langle\cdot,\cdot\rangle$ denotes the inner product. Given an arbitrary function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})$, $\nabla_{\boldsymbol{x}}\boldsymbol{f}(\boldsymbol{x}_{k})\coloneqq\frac{\partial\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})}{\partial\boldsymbol{x}}\Bigr|_{\begin{subarray}{c}\boldsymbol{x}_{k}\end{subarray}}$ and $\nabla_{\boldsymbol{x}\boldsymbol{x}}\boldsymbol{f}(\boldsymbol{x}_{k})\coloneqq\frac{\partial^{2}\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})}{\partial\boldsymbol{x}\partial\boldsymbol{x}}\Bigr|_{\begin{subarray}{c}\boldsymbol{x}_{k}\end{subarray}}$ denote the first-order and second-order partial derivative of $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})$ with respect to $\boldsymbol{x}$ evaluated at $\boldsymbol{x}_{k}$ respectively, and $\nabla_{\boldsymbol{x}\boldsymbol{y}}\boldsymbol{f}(\boldsymbol{x}_{k},\boldsymbol{y}_{k})\coloneqq\frac{\partial^{2}\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})}{\partial\boldsymbol{x}\partial\boldsymbol{y}}\Bigr|_{\begin{subarray}{c}\boldsymbol{x}_{k},\boldsymbol{y}_{k}\end{subarray}}$ denotes the second-order derivative of $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})$ evaluated at $(\boldsymbol{x}_{k},\boldsymbol{y}_{k})$.
2 The Problem
@@
$\boldsymbol{x}(t+1)=\boldsymbol{f}(\boldsymbol{x}(t),\boldsymbol{u}(t)),$ (1)
where $t=0,1,2,\cdots$ denotes the time index, $\boldsymbol{x}(t)\in\mathcal{X}\subset\mathbb{R}^{n}$ and $\boldsymbol{u}(t)\in\mathcal{U}\subset\mathbb{R}^{m}$ denote the system state and control input at time $t$, respectively, and $\boldsymbol{f}:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$ denotes the time-invariant, Lipschitz continuous, and unknown dynamics mapping. We assume $\mathcal{X}$ is a countably infinite state space. Notably, for the remainder of this paper, we distinguish between system state-input variables and fixed observed state-input pairs by denoting $\boldsymbol{x}_{t}$ and $\boldsymbol{u}_{t}$ as constant state and input vectors observed at time $t$, respectively.
@@
$J_{t^{*}}=\sum_{t=t^{*}}^{\infty}\gamma^{t-t^{*}}c(\boldsymbol{x}(t),\boldsymbol{u}(t)),$ (2)
where $c:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$ denotes a bounded and Lipschitz continuous stage cost function, $0<\gamma<1$ is a discount factor, and the control inputs are assumed to be chosen from a policy obeying the following form:
@@
Assume that the closed-loop optimal controller that minimizes $J_{t^{*}}$ in (2) can be represented by $\boldsymbol{\mu}(\boldsymbol{x}(t),\boldsymbol{\theta}^{\mu*})$. The problem of interest is to develop a data-driven framework to achieve $\boldsymbol{\theta}^{\mu*}$. Specifically, given an arbitrary initial state $\boldsymbol{x}(t^{*})=\boldsymbol{x}_{t^{*}}$ at time $t^{*}$, this paper aims to solve the following optimization problem without knowing the actual dynamics $\boldsymbol{f}$ in (1):
$\boldsymbol{\theta}^{\mu*}$ $=\arg\min_{\boldsymbol{\theta}^{\mu}\in\mathbb{R}^{q}}J_{t^{*}}(\boldsymbol{\theta}^{\mu})$ (4)
$\quad\text{subject to:}\quad\boldsymbol{x}(t+1)$ $=\boldsymbol{f}(\boldsymbol{x}(t),\boldsymbol{\mu}(\boldsymbol{x}(t),\boldsymbol{\theta}^{\mu})),\boldsymbol{f}\ \text{unknown},$
$\boldsymbol{x}(t^{*})$ $=\boldsymbol{x}_{t^{*}},\quad\boldsymbol{x}_{t^{*}}\ \text{given}.$
@@
The optimization problem in (4) is a well-established optimal control problem when the dynamics $\boldsymbol{f}$ in (1) is known [lehtomaki1981robustness]. In the case where $\boldsymbol{f}$ is unknown, one approach to solve (4) involves model-free methods, such as RL techniques, which typically require a substantial amount of data to achieve an optimal solution. In the following section, we will propose a framework that concurrently approximates $\boldsymbol{f}$ using the Koopman operator and optimizes $\boldsymbol{\theta}^{\mu}$ to find the $\boldsymbol{\theta}^{\mu*}$. The method is shown to have better efficiency of computing $\boldsymbol{\theta}^{\mu*}$ in later simulations.
@@ 3.1 Challenges and Key Ideas
An iterative method to find $\boldsymbol{\theta}^{\mu*}$ in (4) is to employ the gradient descent method. Specifically, let $k=0,1,2,\cdots$ denote the iteration index, $\boldsymbol{\theta}_{k}^{\mu}\in\mathbb{R}^{q}$ denote the estimation of $\boldsymbol{\theta}^{\mu*}$ at the $k$-th iteration, and $\alpha_{k}^{\mu}$ represent the step size at the $k$-th iteration corresponding to $\boldsymbol{\theta}_{k}^{\mu}$. Given any arbitrary initial $\boldsymbol{\theta}_{0}^{\mu}$, the parameter $\boldsymbol{\theta}_{k}^{\mu}$ is updated iteratively using the gradient of $J_{t^{*}}$ as follows:
$\boldsymbol{\theta}_{k+1}^{\mu}=\boldsymbol{\theta}_{k}^{\mu}-\alpha_{k}^{\mu}\nabla_{\boldsymbol{\theta}^{\mu}}J_{t^{*}}(\boldsymbol{\theta}_{k}^{\mu}),\quad\boldsymbol{\theta}_{0}^{\mu}\ \text{given},$ (5)
@@
$\nabla_{\boldsymbol{\theta}^{\mu}}J_{t^{*}}=\frac{\partial c_{t^{*}}}{\partial\boldsymbol{u}_{t^{*}}}\frac{\partial\boldsymbol{u}_{t^{*}}}{\partial\boldsymbol{\theta}^{\mu}}+\sum_{t=t^{*}+1}^{\infty}\gamma^{t-t^{*}}(\frac{\partial c_{t}}{\partial\boldsymbol{u}(t)}\frac{\partial\boldsymbol{u}(t)}{\partial\boldsymbol{\theta}^{\mu}}$ (6)
@@
Here, $c_{t}\coloneqq c(\boldsymbol{x}(t),\boldsymbol{u}(t))$ is introduced for notation brevity, and the selection of $\alpha_{k}^{\mu}$ will be discussed in detail later in this paper. Throughout this paper, we refer to $\nabla_{\boldsymbol{\theta}^{\mu}}J_{t^{*}}$ as the policy gradient.
Two key challenges arise when computing the policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}J_{t^{*}}$ in (6) due to the unknown dynamics $f$ in (1). First, the system dynamics $\boldsymbol{f}$ is assumed to be unknown in the present work, making the gradient $\frac{\partial\boldsymbol{x}(t)}{\partial\boldsymbol{u}(t-1)}$ unknown as well. Second, calculating $\sum_{t=t^{*}+1}^{\infty}\frac{\partial c_{t}}{\partial\boldsymbol{x}(t)}\frac{\partial\boldsymbol{x}(t)}{\partial\boldsymbol{u}(t-1)}\frac{\partial\boldsymbol{u}(t-1)}{\partial\boldsymbol{\theta}^{\mu}}+\frac{\partial c_{t}}{\partial\boldsymbol{u}(t)}\frac{\partial\boldsymbol{u}(t)}{\partial\boldsymbol{\theta}^{\mu}}$ requires future system states $\boldsymbol{x}(t^{*}+1),\boldsymbol{x}(t^{*}+2),\cdots$ that evolve according to system dynamics. Even if one can approximate the dynamics $\boldsymbol{f}$ with small estimation errors, these errors may still accumulate over an infinite time horizon, potentially resulting in inaccuracies in calculating $\nabla_{\boldsymbol{\theta}^{\mu}}J_{t^{*}}$.
To address the two challenges in computing the policy gradient, we propose a framework that applies the deep Koopman operator (DKO) to approximate the unknown dynamics $\boldsymbol{f}$ in a linear form, enabling efficient approximation of $\frac{\partial\boldsymbol{x}(t+1)}{\partial\boldsymbol{u}(t)}$ and potentially improving gradient-based optimization. Subsequently, we use the temporal difference error technique [sutton1988learning] to approximate $J_{t^{*}+1}$ such that $\nabla_{\boldsymbol{\theta}^{\mu}}J_{t^{*}+1}$ can be estimated using only a one-time-step prediction from the DKO to reduce the computational complexity.